OpenStudy (anonymous):

let n be a positive integer and a,b be invertible integers modulo n such that a is congruent to b^-1 (mod n). What is the remainder when ab is divided by n?

ganeshie8 (ganeshie8):

\[a\equiv b^{-1}\pmod{n}\] multiply both sides by \(b\) and obtain \[ba \equiv bb^{-1}\pmod{n}\] \[ba \equiv 1\pmod{n}\] that means the remainder when \(ab\) is divided by \(n\) is \(1\)

OpenStudy (anonymous):

plain and simple.. this was fast, thanks ganeshie8

OpenStudy (mathmath333):

invertible means inverse ?

ganeshie8 (ganeshie8):

Yes. for example, \(2\) is not invertible in mod \(6\) because there is NO integer satisfying \(2x\equiv 1 \pmod{6}\)

OpenStudy (mathmath333):

ok

OpenStudy (mathmath333):

i am having an observation that \(\large \color{black}{\begin{align} (n-1)^{-1}\pmod n\equiv (n-1) ,n\in \mathbb{N}\hspace{1.5em}\\~\\ \end{align}}\)

ganeshie8 (ganeshie8):

thats right! \[(n-1)(n-1) \equiv (0-1)(0-1) \equiv 1 \pmod{n}\]

ganeshie8 (ganeshie8):

the inverse of `n-1` is `n-1` is same as saying the inverse of `-1` is `-1`

ganeshie8 (ganeshie8):

thats true in any modulo..

OpenStudy (mathmath333):

u mean \(\large \color{black}{\begin{align} (n)^{-1}\pmod m\equiv n ,\quad n<m,\{n,m\}\in \mathbb{N}\hspace{1.5em}\\~\\ \end{align}}\)

ganeshie8 (ganeshie8):

the inverse of `-1` is `-1` in any modulo

ganeshie8 (ganeshie8):

thats true because (-1)(-1) = 1

ganeshie8 (ganeshie8):

Basically inverse of \(a\) in modulo \(n\) is an integer \(x\) which satisfies below congruence \[ax\equiv 1\pmod{n}\]

ganeshie8 (ganeshie8):

when \(a=-1\), we see that \(x=-1\) satisfies the above congruence. so \((-1)^{-1}= -1\)

ganeshie8 (ganeshie8):

this is exact same as our normal multiplcation inverse. the multiplicative inverse of 2 is 1/2 because 2*1/2 = 1

ganeshie8 (ganeshie8):

maybe consider all the inverses in modulo 7 : \[ax\equiv 1\pmod{5}\] \(1^{-1} = 1\) \(2^{-1} = 3\) \(3^{-1} = 2\) \(4^{-1} = 4\)

ganeshie8 (ganeshie8):

\((1)(1) \equiv \color{blue}{1} \pmod{5}\) \((2)(3) \equiv 6\equiv \color{blue}{1} \pmod{5}\) \((3)(2) \equiv 6\equiv \color{blue}{1} \pmod{5}\) \((4)(4) \equiv 16\equiv \color{blue}{1} \pmod{5}\)

ganeshie8 (ganeshie8):

*inverses in modulo \(\color{red}{5}\)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} a\equiv b^{-1}\pmod{n} \end{align}}\) when you multiplied the above by b , then should'nt it be this \(\downarrow\) \(\large \color{black}{\begin{align} ab\equiv b^{-1}b\pmod{nb} \\~\\ ab\equiv 1\pmod{nb} \\~\\ \end{align}}\)

OpenStudy (mathmath333):

@ganeshie8

ganeshie8 (ganeshie8):

yes both are same : \(ab\equiv ba \pmod{n}\)

OpenStudy (mathmath333):

i mean this one \(\large \color{black}{\begin{align} \color{blue}{1.}ab\equiv 1\pmod{nb} \\~\\ \end{align}}\) \(\large \color{black}{\begin{align} \color{blue}{2.}ab\equiv 1\pmod{n} \\~\\ \end{align}}\)

ganeshie8 (ganeshie8):

Ahh okay we have this congruence property : \[a\equiv b \pmod{n} \implies ac\equiv bc\pmod{n}\]

ganeshie8 (ganeshie8):

thats true because \(n|(a-b) \implies n|(ac-bc)\)